A subject has normal values of 125/80 mm Hg blood pressure and heart rate of 65 bpm and a cardiac

output of 5 L min-1.

C. If the heart rate does not change, what is the new peripheral resistance?
Answer #1

The total peripheral resistance can be calculated from the formula

TPR = MAP/ CO

Where TPR is the total peripheral resistance, MAP is the mean arterial pressure and CO is the cardiac output.

Originally the TPR was

TPR = 95 mm Hg / 5 L min-1 = 19.0 mm Hg L-1 min

The new TPR is given as

TPR = 100 mm Hg / 65 bpm x 102.6 mL/ beat = 100/ 6.669 = 14.99 mm Hg L-1 min